Quantity demanded for tea has increased from 100 to 160 units with an increase in the price
of the coffee powder from Rs. 40 to Rs. 50. Calculate the cross elasticity of demand between
tea and coffee and explain the relationship between the goods.
1
Expert's answer
2019-11-22T10:41:24-0500





Cross price elasticity function

"E_c= \frac{\Delta q_x} {\Delta p_y } \times \frac{p_y} {q_x }"

"E_c=" Cross elasticity of demand

"\Delta q_x=" Change in quantity demand of good "x"

"\Delta p_y=" Change in price of good "y"

"p_y=" Initial price of good "y"

"q_x=" Initial quantity demand of good "x"


If cross price elasticity,

>0 then two goods are substitute

=0 then two goods are independent

<0 then two goods are complements



"E_c= \frac{\Delta q_x} {\Delta p_y } \times \frac{p_y} {q_x }"

"E_c= \frac{60} {10 } \times \frac{40} {100 }"

"\boxed{E_c=2.4}"


Cross price elasticity is 2.4 its >0, Then tea and coffee are substitute